# Rethinking Approximate Gaussian Inference in Classification

### v2

@@ Abstract
In classification tasks, softmax functions are ubiquitously used as output activations to produce predictive probabilities. Such outputs only capture aleatoric uncertainty. To capture epistemic uncertainty, approximate Gaussian inference methods have been proposed. We develop a common formalism to describe such methods, which we view as outputting Gaussian distributions over the logit space. Predictives are then obtained as the expectations of the Gaussian distributions pushed forward through the softmax. However, such softmax Gaussian integrals cannot be solved analytically, and Monte Carlo (MC) approximations can be costly and noisy. We propose to replace the softmax activation by element-wise normCDF or sigmoid, which allows for the accurate sampling-free approximation of predictives. This also enables the approximation of the Gaussian pushforwards by Dirichlet distributions with moment matching. This approach entirely eliminates the runtime and memory overhead associated with MC sampling. We evaluate it combined with several approximate Gaussian inference methods (Laplace, HET, SNGP) on large- and small-scale datasets (ImageNet, CIFAR-100, CIFAR-10), demonstrating improved uncertainty quantification capabilities compared to softmax MC sampling.
@@ 1 Introduction
is the $(C-1)$-dimensional probability simplex. The model $\bm{h}$ outputs the probability of inputs $x\in\mathcal{X}$ belonging to each class, $h_{c}(x)=p(c\mid x)$. To learn such a map to the simplex, $\bm{h}$ is typically defined as a composition
@@
For typical models $\bm{h}$ trained through such approximate maximum likelihood estimation, an output probability vector $\bm{h}(x)\in\Delta^{C-1}$ should be interpreted as an estimate of $p_{\text{gen}}(\bm{y}\mid x)$, the probability under the generative model. Notably, the output probability does not take into account the uncertainty of the model due to the finite nature of the data. In the uncertainty quantification formalism, such models can only estimate aleatoric uncertainty and disregard epistemic uncertainty [15].
Figure 1: Mean KL divergence between the ‘true’ predictive (approximated with a $10000$ sample MC approximation) and different predictive approximations on a synthetic data set. This synthetic dataset consists of 100 logit means $\bm{\mu}$ and standard deviations $\bm{\sigma}$, whose components are i.i.d. uniformly distributed. For a fair comparison we scale the logit dataset for the different activations to match the activations to first order, so that $\mu_{c}\in[-1,1]$, $\sigma_{c}\in[0,1]$ for sigmoid, $\mu_{c}\in[-1/2-\log 2,1/2-\log 2]$, $\sigma_{c}\in[0,1/2]$ for softmax, $\mu_{c}\in[-\sqrt{\pi/8},\sqrt{\pi/8}]$, $\sigma_{c}\in[0,\pi/8]$ for normCDF. The MC approximations are capped at a fixed computational budget of $10000$ class samples, i.e. $\lceil 10000/C\rceil$ samples (see also Section C.1).
The probabilistic way to capture epistemic uncertainty is to require the model to output a second-order distribution (that is ‘a distribution over probability distributions’). Namely, for each input $x\in\mathcal{X}$, the model should output a probability measure over the simplex $\mathrm{h}(x)\in\mathcal{P}(\Delta^{C-1})$, as opposed to a point estimate in the simplex $\bm{h}(x)\in\Delta^{C-1}$ [32, 4].
@@
where $\mathrm{f}$ is learned from the training data, $\mathcal{G}(\mathbb{R}^{C})$ is the set of Gaussian measures on the logit space $\mathbb{R}^{C}$, and $\operatorname{\mathbf{softmax}}_{*}$ pushes forward the Gaussian probability measures through the softmax. For example, Heteroscedastic Classifiers (HET) [5] learn the output means and covariances of $\mathrm{f}$ explicitly with a neural network. In linearised Laplace approximations [8], only the means are learned explicitly with a neural network, while the covariances are obtained post-hoc from that neural network by approximating its parameter posterior distribution with a Gaussian, and then pushing it forward to logit space by linearising the network in the parameters. Last-layer Laplace approximations [16] are a variant of linearised Laplace approximations where only the posterior over the last-layer parameters is Gaussian-approximated, requiring no linearisation of the network. Spectral-Normalized Gaussian processes (SNGP) [20] are neural networks with last layers that are approximately Gaussian processes, through random feature expansions and Laplace approximations.
To obtain predictive probabilities from a distributional classifier as in Eq. 5, we marginalise it out by the measure-theoretic change of variables, using $\mathrm{h}(x)=\operatorname{\mathbf{softmax}}_{*}\mathrm{f}(x)$:
@@
In addition to predictive probabilities, the distributional framework allows for the acquisition of other quantities of interest, such as variances, entropy, and information-theoretic decompositions of the predictive uncertainty into aleatoric and epistemic parts [10, 28, 39]. These are usually obtained by Monte Carlo (MC) sampling from the Gaussian measure $\mathrm{f}(x)$. Indeed, while the probability density of $\mathrm{h}(x)=\operatorname{\mathbf{softmax}}_{*}\mathrm{f}(x)$ can be obtained in closed form using a change of variables formula [1], this does not seem of practical use—even predictives cannot be computed analytically. On the other hand, given a variance, the runtime and memory cost required for an MC approximation to achieve an estimator of said variance grows linearly with the number of classes. Crucially, this computational cost is not limited to training time, but also to inference time, when the computational budget may be much more limited. This limits the application of such Gaussian inference methods in classification with a large number of classes, such as those common in computer vision and language modeling.
1.1 Proposed Recipe
We summarise our framework as a 3-step recipe:
@@ 1.1 Proposed Recipe
    Choose an approximate Gaussian inference method that outputs logit means and covariances.
@@
    Choose an output activation that yields a closed-form Gaussian integral (Section 2).
@@
    At inference time, use our closed-form formulas (see Section 2 for the predictives, Section 3 for other quantities of interest).
@@ 2.1 Motivation from Softmax Models
A key difficulty in obtaining exact or closed-form approximate predictive probabilities of $\mathrm{h}(x)=\operatorname{\mathbf{softmax}}_{*}\mathrm{f}(x)$ is that pushforwards of Gaussian distributions through the softmax are intractable. The softmax function is the composition $\bm{\operatorname{\mathbf{softmax}}}=\bm{n}\circ\bm{\exp}$, where $\bm{n}$ is the normalisation function given by $\bm{n}(\bm{q}):=\bm{q}/(q_{1}+\dots+q_{C})$ and $\bm{\exp}$ is applied element-wise. Thus, $\mathrm{h}$ can be written as the composition (see Eq. 5)
@@
Now it is straightforward to see that (Section A.1)
@@ 2.2 The General Framework
The previous argumentation can be applied to a general family of output activation functions: suppose the output activation is of the form $\bm{n}\circ\bm{\varphi}$, where $\bm{\varphi}$ is the element-wise application of some activation function $\varphi\colon\mathbb{R}\to(0,\infty)$. Then we have (see Eq. 7)
@@
Solving the one-dimensional integrals $\int_{\mathbb{R}}\varphi(y)\;\mathcal{N}(\mu,\sigma^{2})(dy)$ (see Eq. 8) allows us to obtain the approximate predictive (see Eq. 10)
@@
For instance, taking $\varphi=\Phi$, the standard Normal cumulative distribution function (normCDF) [^1] The resulting model is distinct from a multinomial probit model, see Appendix B., we have by a classical result (Section A.2)
@@
Taking instead $\varphi=\rho$, the logistic sigmoid, we approximate Eq. 13 using the probit approximation (Section A.3),
@@ 2.3 Theoretical Analysis of the Approximation’s Quality
In Fig. 1, we see empirically that the quality of our predictive approximations outperforms existing predictive approximations on synthetic data. We now explore the quality of our predictive approximations theoretically. We conduct the analysis for the case $\varphi=\rho$. Let
@@
where $\hat{q}$ is the probit approximation (Section A.3). Moreover we write $q_{c}:=q(\mu_{c},\sigma_{c}^{2})$, $\hat{q}_{c}:=\hat{q}(\mu_{c},\sigma_{c}^{2})$ where $(\mu_{c},\sigma_{c}^{2})$ is the logit mean-variance pair for class $c$, $\bm{p}$ for the true predictive and $\bm{\hat{p}}$ for our approximate predictive, i.e.
$\bm{p}:=\mathbb{E}_{\bm{P}\sim(\bm{n}\circ\bm{\rho})_{*}\mathcal{N}(\bm{\mu},\bm{\operatorname{\mathbf{diag}}}(\bm{\sigma}^{2}))}[\bm{P}],\quad\bm{\hat{p}}:=\frac{\hat{q}_{c}}{\sum_{c^{\prime}=1}^{C}\hat{q}_{c^{\prime}}}.$ (18)
To analyse the quality of the approximation, we thus want to bound $\operatorname{D_{KL}}(\bm{p},\bm{\hat{p}})$. Clearly, when all the $Q_{c}$ are deterministic (i.e. constant random variables), where $\bm{Q}\sim\bm{\varphi}_{*}\mathrm{f}(x)$, we have $\bm{p}=\bm{\hat{p}}$. Alternatively, from the derivation Eq. 10, we see that in the limiting case Eq. 12 is exact when $\sum_{c=1}^{C}Q_{c}$ is deterministic. This is a weaker assumption than requiring all $Q_{c}$ to be deterministic individually. In fact, we can derive a rate of decay of $\operatorname{D_{KL}}(\bm{p},\bm{\hat{p}})$ as a function of $\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)$:
@@
Section C.2 provides a formula for $M(\mathcal{K})$ and a proof. We also conduct an empirical investigation of whether the assumptions made in the theorem hold in practice, namely whether the logit means and variances lie in a compact set, and whether $M(\mathcal{K})<\infty$.
The $O$ term in Eq. 19 is due to the splitting of the expectation into the numerator and denominator in the derivation Eq. 10, while $M(\mathcal{K})$ results from the error due to the probit approximation. Importantly, Theorem 2.1 states that the latter does not increase with the number of classes.
@@ 3.1 Dirichlet Matching
The formulation of Eq. 11 does not yet allow this. However, we will now see that one can obtain good Dirichlet approximations by constructing a tractable approximation of Eq. 11 of the form
@@
where $\mathcal{D}(\Delta^{C-1})$ is the space of Dirichlet distributions on $\Delta^{C-1}$. The family of Dirichlets presents an ideal choice of distributions over the simplex as, for such distributions, all quantities of interest are closed-form (see Appendix F). The map $\mathrm{a}$ is constructed by moment matching: we match the first two moments of the Dirichlet distribution to the moments of the Gaussian pushforward $\bm{\varphi}_{*}\mathrm{f}(x)$.
Conveniently, like the first moments described in Section 2, we can obtain second moments for $\exp$
@@
where $T$ is Owen’s T function. For $\rho$, we have the following approximation (Daunizeau [7, Equation 23])
@@
As in Eq. 12, the approximation in Eq. 24 is exact when $\sum_{c=1}^{C}Q_{c}$ is a constant random variable.
If we were to directly attempt to infer the parameters of the Dirichlet $\mathrm{h}(x)$ from Eqs. 12 and 24, we would obtain an overparametrised system of equations with $2C$ equations and $C$ unknowns. Thus, we instead use a classical Dirichlet method of moments [25, Eqs. 19 & 23], which reduces these equations, as we will now describe.
@@
for any $1\leq c\leq C$. So, a natural estimate of $\sum_{c=1}^{C}\gamma_{c}$ that uses all $P_{c}$ is the geometric mean of Eq. 25. Since the mean of the Dirichlet is $\bm{\gamma}/\sum_{c=1}^{C}\gamma_{c}$, we obtain using Eq. 25 an expression for the Dirichlet parameters
@@
Hence, using Eqs. 12 and 24, we obtain the computationally feasible Dirichlet parameters
@@
for $\mathrm{h}(x)$, where $S=\max(\sum_{c=1}^{C}\mathbb{E}[Q_{c}],1)$. Taking the maximum with $1$ in the expression for $S$ is an additional approximation to ensure $\bm{\gamma}>\bm{0}$. Note that by performing such moment matching to obtain the Dirichlet parameters in Eq. 27, the mean of this Dirichlet matches precisely the approximate predictive Eq. 12. That is, the approximate predictive of the exact distributional model (Eq. 11) matches the exact predictive of the approximate distributional model (Eq. 20).

### v3

@@ Abstract
In classification tasks, softmax functions are ubiquitously used as output activations to produce predictive probabilities. Such outputs only capture aleatoric uncertainty. To capture epistemic uncertainty, approximate Gaussian inference methods have been proposed. We develop a common formalism to describe such methods, which we view as outputting Gaussian distributions over the logit space. Predictives are then obtained as the expectations of the Gaussian distributions pushed forward through the softmax. However, such softmax Gaussian integrals cannot be solved analytically, and Monte Carlo (MC) approximations can be costly and noisy. We propose to replace the softmax activation by element-wise normCDF or sigmoid, which allows for the accurate sampling-free approximation of predictives. This also enables the approximation of the Gaussian pushforwards by Dirichlet distributions with moment matching. This approach entirely eliminates the runtime and memory overhead associated with MC sampling. We evaluate it combined with several approximate Gaussian inference methods (Laplace, HET, SNGP) on large- and small-scale datasets (ImageNet, CIFAR-100, CIFAR-10), demonstrating improved uncertainty quantification capabilities compared to softmax MC sampling. Our code is available at github.com/bmucsanyi/probit.
@@ 1 Introduction
is the (C1)(C-1)-dimensional probability simplex. The model $\bm{h}$ outputs the probability of inputs $x\in\mathcal{X}$ belonging to each class, $h_{c}(x)=p(c\mid x)$. To learn such a map to the simplex, $\bm{h}$ is typically defined as a composition
@@
For typical models $\bm{h}$ trained through such approximate maximum likelihood estimation, an output probability vector $\bm{h}(x)\in\Delta^{C-1}$ should be interpreted as an estimate of $p_{\text{gen}}(\bm{y}\mid x)$, the probability under the generative model. Notably, the output probability does not take into account the uncertainty of the model due to the finite nature of the data. In the uncertainty quantification formalism, such models can only estimate aleatoric uncertainty and disregard epistemic uncertainty [hullermeier2021aleatoric].
Figure 1: Mean KL divergence between the ‘true’ predictive (approximated with a $10000$ sample MC approximation) and different predictive approximations on a synthetic data set. This synthetic dataset consists of 100 logit means $\bm{\mu}$ and standard deviations $\bm{\sigma}$, whose components are i.i.d. uniformly distributed. For a fair comparison we scale the logit dataset for the different activations to match the activations to first order, so that $\mu_{c}\in[-1,1]$, $\sigma_{c}\in[0,1]$ for sigmoid, $\mu_{c}\in[-1/2-\log 2,1/2-\log 2]$, $\sigma_{c}\in[0,1/2]$ for softmax, $\mu_{c}\in[-\sqrt{\pi/8},\sqrt{\pi/8}]$, $\sigma_{c}\in[0,\pi/8]$ for normCDF. The MC approximations are capped at a fixed computational budget of $10000$ class samples, i.e. $\lceil 10000/C\rceil$ samples (see also Section˜C.1).
The probabilistic way to capture epistemic uncertainty is to require the model to output a second-order distribution (that is ‘a distribution over probability distributions’). Namely, for each input $x\in\mathcal{X}$, the model should output a probability measure over the simplex $\mathrm{h}(x)\in\mathcal{P}(\Delta^{C-1})$, as opposed to a point estimate in the simplex $\bm{h}(x)\in\Delta^{C-1}$ [sensoy2018evidential, charpentier2020posterior].
@@
where $\mathrm{f}$ is learned from the training data, $\mathcal{G}(\mathbb{R}^{C})$ is the set of Gaussian measures on the logit space $\mathbb{R}^{C}$, and $\operatorname{\mathbf{softmax}}_{*}$ pushes forward the Gaussian probability measures through the softmax. For example, Heteroscedastic Classifiers (HET) [collier2021correlated] learn the output means and covariances of $\mathrm{f}$ explicitly with a neural network. In linearised Laplace approximations [daxberger2021laplace], only the means are learned explicitly with a neural network, while the covariances are obtained post-hoc from that neural network by approximating its parameter posterior distribution with a Gaussian, and then pushing it forward to logit space by linearising the network in the parameters. Last-layer Laplace approximations [kristiadi_being_2020] are a variant of linearised Laplace approximations where only the posterior over the last-layer parameters is Gaussian-approximated, requiring no linearisation of the network. Spectral-Normalised Gaussian processes (SNGP) [liu2020simple] are neural networks with last layers that are approximately Gaussian processes, through random feature expansions and Laplace approximations.
To obtain predictive probabilities from a distributional classifier as in Eq.˜5, we marginalise it out by the measure-theoretic change of variables, using $\mathrm{h}(x)=\operatorname{\mathbf{softmax}}_{*}\mathrm{f}(x)$:
@@
In addition to predictive probabilities, the distributional framework allows for the acquisition of other quantities of interest, such as variances, entropy, and information-theoretic decompositions of the predictive uncertainty into aleatoric and epistemic parts [depeweg2018decomposition, mukhoti2023deep, wimmer2023quantifying]. These are usually obtained by Monte Carlo (MC) sampling from the Gaussian measure $\mathrm{f}(x)$. Indeed, while the probability density of $\mathrm{h}(x)=\operatorname{\mathbf{softmax}}_{*}\mathrm{f}(x)$ can be obtained in closed form using a change of variables formula [atchison_logistic-normal_1980], this does not seem of practical use—even predictives cannot be computed analytically. On the other hand, given a variance, the runtime and memory cost required for an MC approximation to achieve an estimator of said variance grows linearly with the number of classes. Crucially, this computational cost is not limited to training time, but also to inference time, when the computational budget may be much more limited. This limits the application of such Gaussian inference methods in classification with a large number of classes, such as those common in computer vision and language modelling.
Summary of Contributions: Our work extends the classical probit approximation for the Gaussian sigmoid integral in binary classification [spiegelhalter_sequential_1990, mackay_evidence_1992] in the following manner:
@@
    Generalising it to the multi-class setting (Sections˜3.1 and 3.2).
@@
    Analysing it theoretically (Section˜3.3).
@@
    Analysing it empirically, and how it outperforms previous approximations (Figs.˜1 and 7).
  4. (iv)
    Generalising it to second-order distributional Dirichlet approximations (Section˜4).
  5. (v)
    Combining it with advantageous training (Sections˜5 and 7).
Our proposed framework can be summarised as a 3-step recipe:
(i) Choose an approximate Gaussian inference method that outputs logit means and covariances. (ii) Choose an output activation that yields a closed-form Gaussian integral (Section˜3). (iii) At inference time, use our closed-form formulas (see Section˜3 for the predictives, Section˜4 for other quantities of interest).
2 Related Work
In uncertainty quantification, methods that output logit Gaussians, which we refer to as approximate Gaussian inference methods, have become one of the dominant paradigms in uncertainty quantification [collier2021correlated, daxberger2021laplace, kristiadi_being_2020, liu2020simple]. In classification tasks, such methods have the downside of requiring sampling at inference to transform Gaussian distributions into predictive probability vectors.
In parallel, prior work has developed ways to propagate probabilistic uncertainty through neural networks [gast_lightweight_2018]. Several works develop analytic propagation of covariances up to the logits [wright_analytic_2024, li_streamlining_2024]; our method complements these by addressing the propagation of the logit-space Gaussian distributions onto the probability simplex. Few works address this last step. The Laplace bridge [pmlr-v180-hobbhahn22a] proposes a map from Gaussians to Dirichlets on the probability simplex. However, this approach has no guarantees on the quality of its predictives, and it severely underperforms Monte Carlo sampling. Other approaches include the mean-field approximation to the Gaussian softmax integral [lu_mean-field_2021] and the approximations of shekhovtsov_feed-forward_2018. In this work, we propose approximate predictives that outperform existing closed-form predictive techniques and provide second-order Dirichlet distributions for other closed-form quantities of interest. Furthermore, this is the first work to provide a theoretical analysis of such an approximation.
Other works have also proposed various activations for classification to replace the softmax. Particularly relevant is [galy-fajou_multi-class_2020], which composes element-wise sigmoids with normalisation and utilises auxiliary-variable augmentation to obtain closed-form ELBO terms during training; however, its predictives require Monte Carlo estimation. Related conjugate/augmentation-based objectives exist for stick-breaking and tree-structured multi-class GPs [linderman_dependent_2015, achituve_gp-tree_2021], and for BNNs, a tempered-likelihood/Dirichlet perspective with a factorised-Gaussian surrogate has been explored [kapoor_uncertainty_2022]. Yet, all these approaches still require sampling or numerical quadrature at inference time. [spiegelhalter_sequential_1990, mackay_evidence_1992, kristiadi_being_2020] use the probit approximation in the binary case, which we generalise to the multi-class setting. [titsias_rc_aueb_one-vs-each_2016] approximates the softmax by providing lower bounds for point estimate training, but replacing the softmax with these bounds during training does not yield a normalised multi-class likelihood. The binary cross-entropy loss with the element-wise sigmoid activation has been reported to outperform cross-entropy in large-scale settings [wightman_resnet_2021, li_bce_2025]. Our work is the first to use this advantageous loss while normalising the output to obtain proper multi-class probabilistic models with closed-form predictives from logit Gaussians.
Evidential deep learning (EDL) approaches [sensoy2018evidential] offer sample-free outputs through Dirichlet distributions but have been shown to be unreliable for uncertainty estimation and disentanglement [NEURIPS2022_bc1d640f, jürgens2024epistemicuncertaintyfaithfullyrepresented, pandey2023learn, mucsányi2024benchmarking]. We focus on making approximate Gaussian inference methods sample-free in a principled and performant manner; however, for completeness, we compare our approach with EDL methods in Section˜K.8.
Finally, multiple authors have explored continuous approximations to categorical distributions, which enable gradient-based optimization for discrete stochastic models through sampling, such as the Gumbel-max trick [maddison_ast_2014, jang_categorical_2017, maddison_concrete_2017, huijben_review_2023]. These methods principally optimise a single categorical distribution, while the present work is concerned with constructing probability densities over the (continuous) simplex of such categorical distributions.
@@ 3.1 Motivation from Softmax Models
A key difficulty in obtaining exact or closed-form approximate predictive probabilities of $\mathrm{h}(x)=\operatorname{\mathbf{softmax}}_{*}\mathrm{f}(x)$ is that pushforwards of Gaussian distributions through the softmax are intractable. The softmax function is the composition $\bm{\operatorname{\mathbf{softmax}}}=\bm{n}\circ\bm{\exp}$, where $\bm{n}$ is the normalisation function given by $\bm{n}(\bm{q}):=\bm{q}/(q_{1}+\dots+q_{C})$ and $\bm{\exp}$ is applied element-wise. Thus, $\mathrm{h}$ can be written as the composition (see Eq.˜5)
@@
Now it is straightforward to see that (Section˜A.1)
@@ 3.2 The General Framework
The previous argumentation can be applied to a general family of output activation functions: suppose the output activation is of the form $\bm{n}\circ\bm{\varphi}$, where $\bm{\varphi}$ is the element-wise application of some activation function $\varphi\colon\mathbb{R}\to(0,\infty)$. Then we have (see Eq.˜7)
@@
Solving the one-dimensional integrals $\int_{\mathbb{R}}\varphi(y)\;\mathcal{N}(\mu,\sigma^{2})(dy)$ (see Eq.˜8) allows us to obtain the approximate predictive (see Eq.˜10)
@@
For instance, taking $\varphi=\Phi$, the standard Normal cumulative distribution function (normCDF) [^1] The resulting model is distinct from a multinomial probit model, see Appendix B., we have by a classical result (Section˜A.2)
@@
Taking instead $\varphi=\rho$, the logistic sigmoid, we approximate Eq.˜13 using the probit approximation (Section˜A.3),
@@ 3.3 Theoretical Analysis of the Approximation’s Quality
In Fig.˜1, we see empirically that the quality of our predictive approximations outperforms existing predictive approximations on synthetic data. We now explore the quality of our predictive approximations theoretically. We conduct the analysis for the case $\varphi=\rho$. Let
@@
where $\hat{q}$ is the probit approximation (Section˜A.3). Moreover we write $q_{c}:=q(\mu_{c},\sigma_{c}^{2})$, $\hat{q}_{c}:=\hat{q}(\mu_{c},\sigma_{c}^{2})$ where $(\mu_{c},\sigma_{c}^{2})$ is the logit mean-variance pair for class $c$, $\bm{p}$ for the true predictive and $\bm{\hat{p}}$ for our approximate predictive, i.e.
$\bm{p}:=\mathbb{E}_{\bm{P}\sim(\bm{n}\circ\bm{\rho})_{*}\mathcal{N}(\bm{\mu},\bm{\Sigma})}[\bm{P}],\quad\bm{\hat{p}}:=\frac{\hat{q}_{c}}{\sum_{c^{\prime}=1}^{C}\hat{q}_{c^{\prime}}},$ (18)
[…]
@@
Section˜C.2 provides a formula for $M(\mathcal{K})$ and a proof. We also conduct an empirical investigation of whether the assumptions made in the theorem hold in practice, namely whether the logit means and variances lie in a compact set, and whether $M(\mathcal{K})<\infty$.
The $O$ term in Eq.˜19 is due to the splitting of the expectation into the numerator and denominator in the derivation Eq.˜10, while $M(\mathcal{K})$ results from the error due to the probit approximation. Importantly, Theorem˜3.1 states that the latter does not increase with the number of classes.
@@ 4.1 Dirichlet Matching
The formulation of Eq.˜11 does not yet allow this. However, we will now see that one can obtain good Dirichlet approximations by constructing a tractable approximation of Eq.˜11 of the form
@@
where $\mathcal{D}(\Delta^{C-1})$ is the space of Dirichlet distributions on $\Delta^{C-1}$. The family of Dirichlets presents an ideal choice of distributions over the simplex as, for such distributions, all quantities of interest are closed-form (see Appendix˜F). The map $\mathrm{a}$ is constructed by moment matching: we match the first two moments of the Dirichlet distribution to the moments of the Gaussian pushforward $\bm{\varphi}_{*}\mathrm{f}(x)$.
Conveniently, like the first moments described in Section˜3, we can obtain second moments for $\exp$
@@
where $T$ is Owen’s T function. For $\rho$, we have the following approximation (daunizeau_semi-analytical_2017)
@@
As in Eq.˜12, the approximation in Eq.˜24 is exact when $\sum_{c=1}^{C}Q_{c}$ is a constant random variable.
If we were to directly attempt to infer the parameters of the Dirichlet $\mathrm{h}(x)$ from Eqs.˜12 and 24, we would obtain an overparametrised system of equations with $2C$ equations and $C$ unknowns. Thus, we instead use a classical Dirichlet method of moments [minka_2000, Eqs. 19 & 23], which reduces these equations, as we will now describe.
@@
for any $1\leq c\leq C$. So, a natural estimate of $\sum_{c=1}^{C}\gamma_{c}$ that uses all $P_{c}$ is the geometric mean of Eq.˜25. Since the mean of the Dirichlet is $\bm{\gamma}/\sum_{c=1}^{C}\gamma_{c}$, we obtain using Eq.˜25 an expression for the Dirichlet parameters
@@
Hence, using Eqs.˜12 and 24, we obtain the computationally feasible Dirichlet parameters
@@
for $\mathrm{h}(x)$, where $S=\max(\sum_{c=1}^{C}\mathbb{E}[Q_{c}],1)$. Taking the maximum with $1$ in the expression for $S$ is an additional approximation to ensure $\bm{\gamma}>\bm{0}$. Note that by performing such moment matching to obtain the Dirichlet parameters in Eq.˜27, the mean of this Dirichlet matches precisely the approximate predictive Eq.˜12. That is, the approximate predictive of the exact distributional model (Eq.˜11) matches the exact predictive of the approximate distributional model (Eq.˜20).